# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ 4.2 Sequence Convergence
4.2.1 Fixed v Varying Lengths of ZBC Strings
@@ 4.2.2 Growing Sequence Analysis
As shown in Figures 3 C and 3 D, a near-perfect correlation between LZW and the Assembly Index is observed across all conditions (patterned and random) growing sequences. Table 1 displays the averaged Spearman correlations between the assembly index (Ai) and other complexity measures (LZW, Entropy Rate, BDM) for patterned sequences of increasing length. LZW and BDM measures are the most robust, consistently converging to a correlation of 1 across all sequence lengths. The average correlation between LZW and Ai was 0.98-0.99. Similarly, Table 2 displays the convergence of all three complexity measures: LZW, entropy rate and BDM to an averaged Spearman correlation of 1 for the random growing sequences shown in Figures 3 A and 3 B.
4.3 Applications to Molecular Classification
In order to check the claims of the AT authors in [5], we will use their data (presented in Source Data as “MS_data_Standard_Curve.csv”) to investigate the linear relation and Pearson correlation between Ai and complexity measures (LZW and our code for Ai of strings) applied to InChI strings of the molecules. We will also apply such a linear and correlation analysis to demonstrate that the LZW of InChI strings of the molecules are also highly correlated to mass spectrometry data (MS2 peak matrices), which was indicated as one of the main advantages of using MA (i.e., Ai applied to molecules). InChI and MS2 spectra of a molecule are simple representations of the molecular structure and its MS fragmentation pattern, respectively.
We will investigate how MA, LZW, Ai, the length of InChI strings and MS2 spectrometry data can be used to cluster Small Molecules and Peptides, following the examples in [5]. Finally, we will apply eight standard clustering algorithms (hierarchical clustering, Gaussian Mixture Models (GMM), KMeans cluster, SVM, Random Forest (RF), KNN, a neural network (FFNN) and Principal Component Analysis) to separate the data relative to Fig.4 in [5] into three categories: Dead, Abiotic and Biological. We will also use the data from Fig.4 in [5] to illustrate why the authors’ model (a simple linear relation) is easily replaced by other models which are also linear relations between any metric and MS2 peak data. In the end, we demonstrate that the separation between molecules produced by living and non-living organisms proposed in [5] can be reproduced with simple metrics, such as the length of the InChI strings of molecules.
Although we will show that molecular complexity is driven by molecular length, with LZW capable of separating the classes of organic versus nonorganic compounds as reported in [12] several years before the assembly index report on the same test (which featured only about a hundred compounds instead of the extensive 15,000+ complete compound database we investigated), we also showed, in the same article, that BDM was capable of a slight improvement over counting only exact copies in data as implemented by LZW algorithms.
[…]
Acknowledgements
@@
Felipe S. Abrahão acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2021$/$14501$-$8$ and $2023$/$05593$-$1$.
@@
Assembly theory undermines not only long-held ideas in algorithmic information theory, but also elements of influential work throughout complexity theory and systems science, such as statistical theories of molecular evolution [46], the nonequilibrium thermodynamics of self-organisation [47, 48], and evolutionary dynamics conceptualised as multi-level teleonomic processes or ecosystem models of multiscale interactions[49]. For instance, Nobel Laureate and complex systems pioneer Herbert Simon argues that complex systems evolve hierarchical and multi-nested patterns or networks through stable intermediate forms, allowing rapid emergence through scaling of emergent mesoscale structures at distinct levels, from simple to complex [50]. According to physicist Freeman Dyson, for early life to arise and evolve, it needed a means of storing and replicating information with a low error rate [51]. This aligns with ideas in algorithmic information theory about the need for mechanisms to compress information and replicate it accurately. Hence, the ideas of systems theorists such as Simon, Dyson, Prigogine, and Schrödinger about complex systems and negentropy-based self-organisation in living systems, which are essentially saying that life can reduce entropy/disorder in local regions, and that these ordered structures must be able to maintain and transmit their organisation in the form of redundancies.
@@
In the early 2010s, a quantitative measure of physical complexity based on the information required for self-assembly, applicable to any structure, including chemical and molecular networks, was introduced in connection to evolution [61]. By quantifying the complexity of molecules and protein complexes, the approach highlighted how symmetry and modularity are favoured in biological self-assembly, and offered metrics for comparing different physical structures. With the same group, we also reported that a measure of algorithmic complexity of graphs correlated equivalent networks, revealing a connection between automorphism size and algorithmic information content, and showing how to classify real-world graphs, including metabolic networks [62] as in [36]. In this other work in the mid 2010s, we also showed how approximations of algorithmic complexity could quantify the self-assembly properties of non-DNA molecular computing in application to an example of synthetic porphyrin molecules (the molecules that give their colour to red blood cells) [63].
@@
More recently, process biologist Jaeger has pointed out that Assembly Theory does not explain or quantify evolution and selection as claimed, and uses these terms too broadly. Jaeger notes that Assembly Theory’s use of metrics like a structure’s self-assembly through recursively combining objects over minimal paths and predicted abundance based on rule weights is not novel, as they are “special cases” of previously established algorithmic complexity measures like Huffman coding [71] (as approximating Shannon entropy). He points out that Assembly Theory repackages these principles of algorithmic probability and complexity schemes in a new context. Due to the multiscalar complexity and goal-directed behaviours exhibited across living and engineered systems, the traditional binary categories of organic and inorganic do not fully capture this continuum or evolutionary (adaptive) spectrum. Emerging fields in synthetic biology and soft matter systems are demonstrating that life comprises diverse, hybrid forms with programmable, self-producing (autopoietic) properties. Further discussion of extrinsic vs intrinsic complexity measures can be also found in [4].
While the strategy of calling Shannon entropy and $\mathbf{LZ}$ compression by another name has proven fruitful for the authors of Assembly Theory, it has done a disservice to science and good scientific practice and communication. It has proven to be very successful in breaking into the review and news cycle systems, conferring on it a sense of breakthrough and novelty. However, all of AT’s results can be reproduced with naive statistical tools without resorting to new theories or indexes that make grandiose claims, as we have shown here and in previous work [2, 4].
The self-assembly of inert and living matter has been widely recognised and understood for centuries, with many assembly processes explained by physical laws and fundamental principles—from galaxies forming due to gravity to animals and organs constructed from a small subset of polymers. While ignorance has sometimes helped spur new discoveries, Assembly Theory offers no new insights into the mechanisms driving these assembly processes and is no more capable of quantification than other algorithms from the existing literature, some of which have been previously reported but overlooked by the authors [12, 22].
@@
On the overreaction against compression: the scientific endeavour is about compression through model simplification, to explain more with less, effectively compressing observed data into scientific theories. In fact, compression has been proven to be equivalent to prediction [27], and even equivalent to comprehension broadly speaking [72], so it is a fundamental concept of science [40], not just of computer science, and not a technicality bearing no relevance to nature. It is a universal assumption in the practice of science based on Ockham’s razor [73]: the better the compression the more likely a model or theory is to capture a natural phenomenon, and the more accurately it may predict its behaviour. This has been a dictum of science since the times of Galileo, is foundational to formal science and is at the core of Western philosophy, and it has nothing to do with the technicalities of compression, which the authors of AT seem to overreact to in their attempt to distance themselves from actual specific algorithms introduced in the 1970s [74], from which theirs happen to be indistinguishable.
The authors of AT have not understood the value of compression as a concept fundamental to science, and how science and the concept of explanation and model are deeply connected to the concept of compression [75]. Indeed, compression and prediction, for example, are formally indistinguishable [39, 41, 76, 38]: if you can compress something you have gained an understanding of the underlying patterns that allow you in principle to predict future outcomes; if you cannot, then prediction is not possible [27]. These connections were made by the founders of algorithmic information theory, showing an equivalence between compression, prediction and universal statistical tests for randomness, assuming nothing else but the concept of computation and the existence of computable processes.
The sort of crude understanding of complexity hierarchies and the relationship with other indexes displayed in e.g. [1, 23] explains some of the field’s shady reputation at times. Authors keep producing indexes that simply turn out to be redundant, without a deep understating of the way in which they can be inserted into the complexity landscape, and neglecting to perform the proper, most basic control experiments and comparisons.
If Assembly Theory is looking for the most likely causal explanations for agnostic selective and evolutionary choices (i.e., without taking into consideration the agent’s interaction with its environment), then it is not copy number but algorithmic probability [9, 34] that it should be looking at and applying [4].
@@
In summary, our findings demonstrate that the authors of Assembly Theory still have to explain, account for, and show how their assembly index is different from and more useful compared to existing indexes, in particular those of the $\mathbf{LZ}$ family (and, as demonstrated in this article, even LZW when used on more than a narrow selection of examples). There has been a traditional complexity science or complex systems community absorbed in producing endless variations of Shannon entropy measures over the years and decades, with no reasonable explanation of how they are inserted into the hierarchy of fundamental measures led by Shannon entropy and algorithmic complexity, beyond simplistic and completely misleading accounts (such as Fig. 1 in [23]). Assembly Theory is one such measure that claims to be orthogonal to Shannon entropy or algorithmic complexity but is not [2, 4]. Whether it has any pedagogical merit remains to be proven by the authors, but it has no fundamental character or methodological value.
We believe that science and mathematics will prevail in exposing measures presented as different or extraordinary when in fact they are derivative.
@@
In order to compare the groups in Fig. 9, we perform a pair-wise Cucconi test [77, 78] to check if each pair can be considered as coming from distributions with the same location and scale. The pair-wise p-values can be found in Fig. 11.
Figure 11: Heatmap showing the p-values for pair-wise Cucconi tests for all the groups presented in Fig. 9. The p-values greater than 0.05 were also explicitly written. It is clear that all the algorithms considered are able to separate between classes, because the pair-wise p-values reject the null hypothesis (equality of location and scale) with 95% confidence. Also, in accordance with our previous claims, Ai and LZW (both for the same group of compounds) were tested for equal location and scale and such a null hypothesis was not rejected. This was conducted on the same small set of molecular compounds picked by the authors [5] unlike the exhaustive experiment we reported years before [12].
The p-values in Fig. 11 can be used to assess if the groups formed, when analysed by the calculated metrics, are statistically similar. For example:
@@
  • For a 95% level of confidence, we can say that the null hypothesis is rejected (p-value $<$ 0.05). Therefore, since the null hypothesis is that both groups have the same location and scale, we can say that the two groups are not statistically similar (location-scale wise, which is much stronger than simply comparing the means).
@@
Table 4: Between group comparison results of various complexity measures from the box plot shown in 9. The normality tests (Shapiro-Wilk, given by P-values and a pass or fail for normality) were followed by the non-parametric Kruskal-Wallis test for comparing the complexity measures (algorithms) between independent samples: peptides and small molecules. All measures show significant differences between the two molecule types. Based on the P-values, MA can distinguish between the two categories of molecules with a performance comparable to the most naive statistical indexes and on all molecular descriptors (InChI length and MS2 spectra). This was conducted on the same small set of molecular compounds picked by the authors [5], unlike the exhaustive experiment we reported years before [12].
Statistically, as shown in Table 4, there is no fundamental difference or advantage to using Ai over traditional methods, or utilising one or another data type, as we have shown that the main driver is simply molecular length and not any other feature, including ‘copy-number’, not to mention all the false negatives that were misclassified by MA and were not highlighted as the ones above the mistaken and arbitrary threshold.
@@
    Assembly Theory would need to demonstrate that it can do something that other multiple stochastic measures cannot. Indeed, most complexity measures are stochastic in nature or by design. One argument could be that the assembly index combines this stochasticity with a mechanical assembly procedure, but it would still need to show it can do something that none of the other statistical measures, such as Shannon entropy, can do, or that it is also different from $\mathbf{LZ}$ algorithms. The evidence so far, however, is that when it comes to the tasks that Assembly Theory was supposed to innovate and pioneer, not only LZW but Shannon entropy and other naive statistical indexes are able to replicate or outperform AT.
@@
Nevertheless, let us briefly entertain the possibility that AT is doing something different than, say $\mathbf{LZ}$ algorithms, because of its ‘stochastic nature’. It should then have an effect on its resulting calculations and produce results different from other non-stochastic approaches. However, this is what we have said all AT papers have lacked, and what is considered in science a basic control experiment, an elementary comparison to previous literature and existing knowledge, which AT has never supplied. However, we have, and the empirical results are that their measure is identical to $\mathbf{LZ}$ compression. When inspected formally, we also demonstrated that their definition of the assembly index is indeed an $\mathbf{LZ}$ grammar that approximates Shannon entropy at basically the same rate as $\mathbf{LZ}$ compression algorithms.
If AT takes into consideration or works in a different way yet correlates with the most popular compression algorithm to date and at the greatest statistical significance, what would the value of any feature attributed to it be? Even assuming such a stochastic feature, AT cannot capture anything beyond the realm of statistical triviality and therefore ends up performing just as any fully deterministic algorithm, as we have proven both empirically and theoretically. Thus it represents nothing fundamentally distinctive, nor does it provide any insights other than perhaps pedagogical or accidental insights. In other words, if AT is Shannon entropy implemented by an $\mathbf{LZ}$ compression algorithm with extra steps, why would researchers want the extra steps?
Figure 12: No Turing machine was harmed: this diagram shows the stochastic and computational components of the Block Decomposition Method applied to any object $s$, finite, distinguishable and breakable (all properties that AT has deemed to innovate in dealing with when faced with having to define an object). $s$ can be any type of object as we have shown before, not only binary strings but matrices, vectors, images, tensors, networks and objects in any vocabulary (binary or not). BDM traverses all the possible pathways from the most likely origin to the reconstruction of an object $s$ block by block, with the additional advantage that it can deal with and account for non-trivial causal regularities accounting for any operation and not only join operations as AT does, operations including but not restricted to deletion, subtraction, reversion, complementation, transposition, rotation, inversion, and so on. CTM, as an optimal inference method based on algorithmic probability, is the other side of the symbolic AI coin with BDM, the neurosymbolic approach that combines global statistical entropic distances with local symbolic algorithmic complexity distances. While there is a process to increase the accuracy of CTM at finding more and better computer programs, and such a process is semicomputable, CTM and BDM themselves are fully computable, and BDM runs in linear time.
@@
In Fig. 12, the sum of all the pathways weighted by algorithmic probability (shortest and fastest first, longer later) on the basis that shorter explanations that can explain more of the object are favoured, which is consistent with the science dictum of model bias that can explain more with less and explored in our own papers(e.g. [75]). BDM can account for all causal operations and not only join operations like AT. See also Section 4.1.

### v10

@@ 4.2 Sequence Convergence
4.2.1 Fixed vs. Varying Lengths of ZBC Strings
@@ 4.2.2 Growing Sequence Analysis
As shown in Figures 2C and 2D, a near-perfect correlation between LZW and the Assembly Index is observed across all conditions (patterned and random) growing sequences. Table 1 displays the averaged Spearman correlations between the assembly index (Ai) and other complexity measures (LZW, Entropy Rate, BDM) for patterned sequences of increasing length. LZW and BDM measures are the most robust, consistently converging to a correlation of 1 across all sequence lengths. The average correlation between LZW and Ai was 0.98-0.99. Similarly, Table 2 displays the convergence of all three complexity measures: LZW, entropy rate, and BDM to an average Spearman correlation of 1 for random growing sequences shown in Figures 2A and 2B.
5 Conclusions
Table 5: What we have demonstrated based on mathematical and empirical evidence.
Claim by AT/Ai Our counter-argument Technical basis
AT/Ai explain selection/evolution Ai measures redundancy, not environment-dependent selection Selection depends on $(o,e)$; Ai $\sim$ LZ cannot encode fitness or counterfactuals
AT unifies physics and biology No new invariant/law; Ai reduces to entropy-like redundancy Representation-dependent score; no conserved quantity or predictive law
Ai is not data compression Ai is a dictionary parsing (statistical compression) algorithm proven mathematically with an undisputed theorem [17] Equivalent to LZ-style factorisation; code-length $\propto$ Ai output [5, 17]
Ai is unrelated to Shannon or Kolmogorov complexity Ai approximates expected code length; loosely upper-bounds $K(x)$ Hierarchy: Ai $\equiv$ LZ $\preceq H\prec$ BDM $\preceq$ CTM $\prec$ AP $\equiv K$
Kolmogorov complexity is unusable because it is uncomputable $K$ is semi-computable; many branches of mathematics rely on uncomputable concepts (e.g., calculus, analysis) BDM/CTM give computable $K$-proxies with many applications [39, 40]
AT avoids Turing-machine limits Ai is computed on a computer program, hence a Turing machine; appealing to time classes is inconsistent Rejecting TMs while invoking time complexity is self-contradictory [4]
Compression destroys physical meaning Compression units or tokens can represent molecules rather than binary strings; the objection confuses algorithm with dictionary particularities Fixing dictionary units to molecular granularity keeps objects intact but even ‘destroying’ them, they are recovered from their patterns
Ai uniquely captures stochastic/causal structure Algorithmic probability and Shannon entropy already handle uncertainty; Ai adds extra steps and unfounded claims Shannon entropy, Kolmogorov complexity, and algorithmic probability subsume copy-counting [17]
Any positive result reported by AT is not the consequence of a unique breakthrough or feature of AT but the direct result of computing Shannon entropy under another name The same outcomes, including cutoff values distinguishing organic from inorganic compounds, have been previously reported in the literature using entropy and compression-based analyses, as well as methods that go beyond counting trivial repetitions and justify causality [22] AT reproduces prior entropy-derived results and presents them as novel while offering no new theoretical or empirical contribution [22, 29]
Table 6: What we have neither claimed nor advocated for.
Not our claim Our position / clarification
Statistical compressors should replace AT as a theory of selection and evolution Compression serves as a null benchmark revealing Ai’s redundancy, not a theory of life. AT and Ai implement statistical compression and is as inadequate to characterise evolution and selection as any other statistical compression algorithm
Ai is inefficient and should be swapped for faster compressors Our critique is conceptual and foundational; runtime is not the issue. It is the AT authors who claim that because their algorithm is slower than statistical compressors, it is different and better with no evidence
Molecules are just strings The application of data compression to e.g. SMILES molecular descriptions in our experiments had the only purpose to serve as a control experiment to show that it produces the same results that AT claims is capable of with Ai on experimental data
Shannon entropy or Kolmogorov complexity can explain selection or evolution We do not reductionistically derive biology from either measure; we show that AT adds no new insight, whereas Kolmogorov complexity can in fact capture causality unlike AT
Entropy should replace AT as a theory We advocate removing methodological redundancy, not elevating $H$ to a theory of life
Algorithmic similarity implies ontological identity Informational equivalence shows lack of novelty, not sameness of phenomena
Given that our position has been oversimplified by some authors, who have implied that we advocate the use of compression algorithms or their application to SMILES as an alternative to Assembly Theory (AT) [41], we have clarified and summarised our position in Tables 5 and 6. In fact, our argument is the opposite: we have shown that AT, when implemented with Ai, is effectively a trivial statistical compression algorithm. As such, it cannot account for selection or evolutionary processes any better than other algorithms in the same class. Moreover, applying any statistical compression algorithm to SMILES molecular representations reproduces the same results reported by AT, thereby casting doubt on its novelty or the fundamental nature of its contribution.
As we have shown, it is not possible for Assembly Theory (AT) and its Assembly Index (Ai) to explain or quantify selection or evolution because of its trivial nature, which does not distinguish it from Shannon entropy, neither fundamentally nor judging by its results. Instead, AT’s sophisticated jargon collapses the distinction between causal and descriptive domains, confusing the historical processes that generate diversity with the static configurations that these processes leave behind. Contrary to their claims, AT takes objects as instant and final, for which they produce plausible statistical, not causal, origins with no evidence other than chemical or molecular constructibility completely disconnected from evolution and selection as Ai only captures the minimal number of trivial combinatorial steps required to build a molecule, rather than the causal or heritable dynamics through which biological change unfolds. This makes AT fundamentally incapable of representing evolutionary mechanisms such as variation, inheritance, and selection, which depend on temporally ordered transformations rather than structural enumeration.
The assumption that statistical repetition (molecular or otherwise) mirrors evolutionary progression fails in reducing all differences to a superficial similarity metric. In essence, AT substitutes pattern resemblance for causal explanation, offering a phenomenological index of statistical difficulty that Shannon entropy already captures but without accounting for how or why those statistical regularities may accumulate over time. AT confounds speculation with theory.
This paper completes the theoretical and empirical demonstration that any variation of the copy-number concept underlying AT and Ai, which involves counting the number of repetitions in an object, such as a molecule, to arrive at a measure of selection or evolution, is equivalent to popular statistical compression algorithms based on Shannon entropy. We showed that the authors’ argument that the intractability of AT separates Ai from standard compression does not withstand basic scrutiny and is not robust, and that their empirical results separating organic from inorganic compounds, on which their whole empirical evidence is based, have not only been previously reported, avoiding claims of grand unification between physics and biology, but are mostly driven by molecular length—which the authors did not control for.
Given that Ai can only account for basic statistical operations that cannot account for the full extent and complexity of selection and evolution, this contribution makes AT unsuitable and redundant, especially when applied to their own experimental data, on which we have proven that both Shannon entropy and LZ compression reproduce the same or even better results without the extra steps and the additional unjustified assembly argot that sometimes can be useful if honest, open and transparent.
What AT has offered is an obscure theory-of-everything narrative with little substance and a refusal to engage with control comparisons or foundational mathematics. At best, we consider Assembly Theory (AT) a pedagogical exercise, potentially useful for illustrating certain patterns of structural repetition in molecular complexity that might or might not lead to future insights. However, this limited didactic value stands in clear contrast to the sweeping claims made by its proponents in high-profile media coverage, exaggerated press releases, and academic publications. AT does not explain evolution or selection; it does not unify physics and biology; and it does not redefine complexity. As our analyses demonstrate, AT simply reframes familiar statistical compression methods in unfamiliar terminology, introduces unnecessary computational steps, and presents the result as if it were a scientific breakthrough.
6 Funding
@@
Felipe S. Abrahão acknowledges support from the São Paulo Research Foundation (FAPESP), grants 20212021/1450114501-88 and 20232023/0559305593-11. No other funding was received.
@@
Assembly theory has undermined long-held ideas in algorithmic information theory, but also elements of influential work throughout complexity theory and systems science, such as statistical theories of molecular evolution [42], the nonequilibrium thermodynamics of self-organisation [43, 44], and evolutionary dynamics conceptualised as multi-level teleonomic processes or ecosystem models of multiscale interactions[45]. For instance, Nobel Laureate and complex systems pioneer Herbert Simon argues that complex systems evolve hierarchical and multi-nested patterns or networks through stable intermediate forms, allowing rapid emergence through scaling of emergent mesoscale structures at distinct levels, from simple to complex [46]. According to physicist Freeman Dyson, for early life to arise and evolve, it needed a means of storing and replicating information with a low error rate [47]. This aligns with ideas in algorithmic information theory about the need for mechanisms to compress information and replicate it accurately. Hence, the ideas of systems theorists such as Simon, Dyson, Prigogine, and Schrödinger about complex systems and negentropy-based self-organisation in living systems, which are essentially saying that life can reduce entropy/disorder in local regions, and that these ordered structures must be able to maintain and transmit their organisation in the form of redundancies.
@@
In the early 2010s, a quantitative measure of physical complexity based on the information required for self-assembly, applicable to any structure, including chemical and molecular networks, was introduced in connection to evolution [58]. By quantifying the complexity of molecules and protein complexes, the approach highlighted how symmetry and modularity are favoured in biological self-assembly, and offered metrics for comparing different physical structures. With the same group, we also reported that a measure of algorithmic complexity of graphs correlated equivalent networks, revealing a connection between automorphism size and algorithmic information content, and showing how to classify real-world graphs, including metabolic networks [59] as in [12]. In this other work in the mid 2010s, we also showed how approximations of algorithmic complexity could quantify the self-assembly properties of nonDNA molecular computing in application to an example of synthetic porphyrin molecules (the molecules that give their colour to red blood cells) [60].
@@
More recently, process biologist Jaeger has pointed out that Assembly Theory does not explain or quantify evolution and selection as claimed, and uses these terms too broadly. Jaeger notes that Assembly Theory’s use of metrics like a structure’s self-assembly through recursively combining objects over minimal paths and predicted abundance based on rule weights is not novel, as they are “special cases” of previously established metrics [68].
Due to the multiscalar complexity and goal-directed behaviours exhibited across living and engineered systems, the traditional binary categories of organic and inorganic do not fully capture this continuum or evolutionary (adaptive) spectrum. Emerging fields in synthetic biology and soft matter systems are demonstrating that life comprises diverse, hybrid forms with programmable, self-producing (autopoietic) properties. Further discussion of extrinsic vs. intrinsic complexity measures can also be found in [17].
While the strategy of calling Shannon entropy and LZ compression by another name has proven fruitful for the authors of Assembly Theory, it has done a disservice to science and good scientific practice and communication. It has proven to be very successful in breaking into the review and news cycle systems, conferring on it a sense of breakthrough and novelty. However, all of AT’s results can be reproduced with naive statistical metrics without resorting to new theories or indexes that make grandiose claims, as we have shown here and in previous work [5, 17].
The self-assembly of inert and living matter has been widely recognised and understood for centuries, with many assembly processes explained by physical laws and fundamental principles—from galaxies forming due to gravity to animals and organs constructed from a small subset of polymers. Although ignorance has sometimes helped spur new discoveries, Assembly Theory does not offer new insights into the mechanisms driving these assembly processes and is not more capable of quantification than other algorithms in the existing literature, some of which have been previously reported but have been overlooked by the authors [22, 29].
@@
On the overreaction against compression: the scientific endeavour is about compression through model simplification, to explain more with less, effectively compressing observed data into scientific theories. In fact, compression has been shown to be equivalent to prediction [21], and even equivalent to comprehension, broadly speaking [69], so it is a fundamental concept of science [70], not just computer science, and not a technicality bearing no relevance to nature. It is a universal assumption in the practice of science based on Ockham’s razor [71]: the better the compression, the more likely a model or theory is to capture a natural phenomenon and the more accurately it may predict its behaviour. This has been a dictum of science since the times of Galileo, is foundational to formal science, and is at the core of Western philosophy, and it has nothing to do with the technicalities of compression, which the authors of AT seem to overreact to in their attempt to distance themselves from actual specific algorithms introduced in the 1970s [9], from which theirs happen to be indistinguishable.
The authors of AT appear not to realise the value of compression as a fundamental concept and practice in science, and how explanation and model are deeply connected to compression [72] both statistically and algorithmically. Compression and prediction have been formally connected [14, 73, 74, 10] showing that if one can compress an object, one has gained an understanding of the underlying patterns that allow prediction; if not, then prediction is impossible [21]. These connections were made by the founders of algorithmic information theory, showing an equivalence between compression, prediction and universal statistical tests for randomness, assuming nothing else but the concept of computation and the existence of computable processes.
The sort of crude understanding of complexity hierarchies and the relationship with other indexes displayed in e.g. [4, 26] explains some of the field’s shady reputation at times. Authors keep producing indexes that simply turn out to be redundant, without a deep understating of the way in which they can be inserted into the complexity landscape, and by neglecting to perform the most basic control experiments and comparisons, new metrics can make a disservice to the field, especially when introduced to the public together with unfounded claims, such as the claims that AT unifies biology and physics by way of simply counting repetitions in data.
If Assembly Theory is looking for the most likely causal explanation for agnostic selective and evolutionary choices (i.e., without taking into consideration the agent’s interaction with its environment), then it is not copy number but algorithmic probability [8, 28] that AT should be studying and applying [17].
@@
In practice, all real-world applications of Kolmogorov complexity involve computable approximations, such as LZW, CTM, or BDM, each progressively approaching the uncomputable ideal. Ai, being computable and empirically defined, belongs squarely among these approximations rather than transcending them.
Our findings demonstrate that the authors of Assembly Theory still have to explain, account for, and show how their assembly index is different from—and more useful to— existing indexes, in particular those of the LZ family (and, as demonstrated in this article, even LZW when used on more than a narrow selection of examples). There has been a traditional complexity science or complex systems community absorbed in producing endless variations of Shannon entropy measures over the years and decades, with no reasonable explanation of how they are inserted into the hierarchy of fundamental measures led by Shannon entropy and algorithmic complexity, beyond simplistic and completely misleading accounts (such as Fig. 1 in [26]).
@@
In order to compare the groups in Fig. 6, we perform a pairwise Cucconi test [75, 76] to check if each pair can be considered to come from distributions with the same location and scale. The pair-wise p-values can be found in Fig. S1.
Figure S1: Heatmap showing the p-values for pair-wise Cucconi tests for all the groups presented in Fig. 6. The p-values greater than 0.05 were also explicitly written. It is clear that all the algorithms considered are able to separate between classes, because the pair-wise p-values reject the null hypothesis (equality of location and scale) with 95% confidence. Also, in accordance with our previous claims, Ai and LZW (both for the same group of compounds) were tested for equal location and scale and such a null hypothesis was not rejected. This was conducted on the same small set of molecular compounds picked by the authors [1] unlike the exhaustive experiment we reported years before [22].
The p-values in Fig. S1 can be used to assess if the groups formed, when analysed by the calculated metrics, are statistically similar. For example:
@@
  • For a 95% level of confidence, we can say that the null hypothesis is rejected (p-value << 0.05). Therefore, since the null hypothesis is that both groups have the same location and scale, we can say that the two groups are not statistically similar (location-scale wise, which is much stronger than simply comparing the means).
@@
Table S1: Between group comparison results of various complexity measures from the box plot shown in 6. The normality tests (Shapiro-Wilk, given by P-values and a pass or fail for normality) were followed by the nonparametric Kruskal-Wallis test for comparing the complexity measures (algorithms) between independent samples: peptides and small molecules. All measures show significant differences between the two molecule types. Based on the P-values, MA can distinguish between the two categories of molecules with a performance comparable to the most naive statistical indexes and on all molecular descriptors (InChI length and MS2 spectra). This was conducted on the same small set of molecular compounds picked by the authors [1], unlike the exhaustive experiment we reported years before [22].
Statistically, as shown in Table S1, there is no fundamental difference or advantage to using Ai over traditional methods, or utilising one or another data type, as we have shown that the main driver is simply molecular length and not any other feature, including ‘copy-number’, not to mention all the false negatives that were misclassified by MA and were not highlighted as the ones above the mistaken and arbitrary threshold.
@@
    Assembly Theory would need to demonstrate that it can do something that other multiple stochastic measures cannot. Indeed, most complexity measures are stochastic in nature or by design. One argument could be that the assembly index combines this stochasticity with a mechanical assembly procedure, but it would still need to show it can do something that none of the other statistical measures, such as Shannon entropy, can do, or that it is also different from LZ algorithms. The evidence so far, however, is that when it comes to the tasks that Assembly Theory was supposed to innovate and pioneer, not only LZW but Shannon entropy and other naive statistical indexes are able to replicate or outperform AT.
@@
Nevertheless, let us briefly entertain the possibility that AT is doing something different than, say LZ algorithms, because of its ‘stochastic nature’. It should then have an effect on its resulting calculations and produce results different from other nonstochastic approaches. However, this is what we have said all AT papers have lacked, and what is considered in science a basic control experiment, an elementary comparison to previous literature and existing knowledge, which AT has never supplied. However, we have, and the empirical results are that their measure is equivalent to LZ compression. When inspected formally, we also demonstrated that their definition of the assembly index is indeed an LZ grammar that approximates Shannon entropy at basically the same rate as LZ compression algorithms.
If AT takes into consideration or works in a different way yet correlates with the most popular compression algorithm to date and at the greatest statistical significance, what would the value of any feature attributed to it be? Even assuming such a stochastic feature, AT cannot capture anything beyond the realm of statistical triviality and therefore ends up performing just as any fully deterministic algorithm, as we have proven both empirically and theoretically. Thus it represents nothing fundamentally distinctive, nor does it provide any insights other than perhaps pedagogical or accidental insights. In other words, if AT is Shannon entropy implemented by an LZ compression algorithm with extra steps, why would researchers want the extra steps?
Figure S2: This diagram shows the similarities of our BDM that predates AT for many years and how its stochastic and computational components applied to any object $s$, finite, distinguishable and breakable (all properties that AT has deemed to innovate in dealing with when faced with having to define an object). $s$ can be any type of object as we have shown before, not only binary strings but matrices, vectors, images, tensors, networks and objects in any vocabulary (binary or not). BDM traverses all the possible pathways from the most likely origin to the reconstruction of an object $s$ block by block, with the additional advantage that it can deal with and account for nontrivial causal regularities accounting for any operation and not only join operations as AT does, operations including but not restricted to deletion, subtraction, reversion, complementation, transposition, rotation, inversion, and so on. CTM, as an optimal inference method based on algorithmic probability, is the other side of the symbolic AI coin with BDM, the neurosymbolic approach that combines global statistical entropic distances with local symbolic algorithmic complexity distances. While there is a process to increase the accuracy of CTM at finding more and better computer programs, and such a process is semicomputable, CTM and BDM themselves are fully computable, and BDM runs in linear time.
@@
In Fig. S2, the sum of all the pathways weighted by algorithmic probability (shortest and fastest first, longer later) on the basis that shorter explanations that can explain more of the object are favoured, which is consistent with the science dictum of model bias that can explain more with less and explored in our own papers(e.g. [72]). BDM can account for all causal operations and not only join operations like AT. See also Section 4.1.